OpenStudy (anonymous):

Which equation is a quadratic equation? 3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3)

OpenStudy (callisto):

Quadratic question: highest degree of polynomial is 2 For (1), 3x-y = 8x+2 y = -5x-2 the highest degree of polynomial is one . For (2) , expand the expression. y = 7x(5 − 2x) + 9 y = 35x - 14x^2 +9 = -14x^2 +35x +9 What is the highest degree of polynomial ? For (3) , expand the expression. 7x^2 + y = (3 + x)(7x − 1) 7x^2 +y = 7x^2 + 21x - x -3 y = 20x -3 What is the highest degree of x ? For (4), expand the expression. y − 3 = (x^2 − 3)(8x + 3) y - 3 = 8x^3 + 3x^2 - 24x -9 y = 8x^3 + 3x^2 - 24x - 6 What is the highest degree of polynomial ? I think you can get the answer now~